OpenStudy (anonymous):

To assess the accuracy of a kitchen scale, a standard weight, known to weigh 1 gram, is weighed a total of n times, and the mean of the n weight measurements is computed. Suppose the scale readings are normally distributed with unknown mean μ and standard deviation σ = 0.01 gram. How large should n be so that a 90% confidence interval for μ has a margin of error of E = ± 0.0001 gram?

OpenStudy (anonymous):

Choices are: A. 165 B. 27,061 C. 38,416

OpenStudy (anonymous):

Margin of error = z-score for 90% confidence*standard deviation/sqrt n 0.0001 = 1.645*0.01/sqrt n sqrt n = 1.645*0.01/0.0001 = 164.5 on squaring both sides n = 27061 Therefore answer to the question is CHOICE (B)...27,061